OpenStudy (anonymous):

(2X-3) (4-3X) Find the product of the binomials. A)-6x^2 + 17x - 12 B)11x^2 - 12 C)17x^2 + 17x - 12 D)72x^2

OpenStudy (raden):

use the FOIL 2x * 4 = 8x -3*4 + 2x * -3x = -12 - 6x^2 -3 * -3x = 9x add up all, then simplify it is 8x - 12 - 6x^2 + 9x = -6x^2 + 17x - 12

OpenStudy (anonymous):

So A?

OpenStudy (raden):

yup :)